Commander Report – December 4th

Sol: 7
Summary Title: EVA #9 and Suit Troubleshooting
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. The crew enjoyed an egg, cheese, and vegetable casserole for breakfast. After a morning planning meeting, HSO collected crew health stats. CDR, Technology Officer, and Green Hab Officer conducted a short EVA in the early afternoon to deploy the ham radio antenna and fly two drone flights around the Hab and around the mirror coating experiment sample location. Fortunately, winds were light today, so the drone was able to successfully complete both flights. During the EVA, Engineer and HSO completed troubleshooting of the two-piece EVA suits and determined that the power strip the charging cables were plugged into had gone bad. They completed further tests with various combinations of suits, batteries, and chargers to ensure all six suits are now charging and holding charge properly. The crew relaxed a bit in the afternoon, reading, knitting, working on art projects, and sharing photos and stories. For dinner, the crew enjoyed tacos with homemade tortillas and baked a cake to celebrate the midpoint of their mission.

Copyright © The Mars Society. All rights reserved. | Main Site